OpenStudy (anonymous):

whats the equation of a line that passes through (-1,3) and (-3,7)??

sam (.sam.):

use slope formula \[slope=\frac{y-y_1}{x-x_1}\]

OpenStudy (anonymous):

haha i know, slope is -2 but everytime i get an answer, it doesnt match the point so im doing something wrong

sam (.sam.):

then pick any ONE of the two points, then use the formula \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

i did! and the equation comes up as y=-2x+11, it doesnt match the point

OpenStudy (anonymous):

when you plug in the numbers

sam (.sam.):

I get y=-2x+1, did you do the subtraction correctly?

sam (.sam.):

y-7=-2(x-(-3)) y=-2x-6+7

OpenStudy (anonymous):

oh, thanks! i mistakenly put the -3 on y and the 7 on x